# MCGMark: An Encodable and Robust Online Watermark for LLM-Generated Malicious Code

### v1

@@ 3.2 Data Pre-processing
To filter out real instances and potential scenarios of malicious code generation by LLMs, the data preprocessing process is divided into two steps, as shown in Fig. 2. First, we filter out extraneous data, such as empty code repositories and duplicate results. Second, we extract instances of malicious code from the filtered results to help us understand the functionality of malicious code and construct the MCGTest dataset.
@@
To ensure the reliability of this analysis. Four researchers with more than four years of software development experience are assigned to filter and analyze the collected results. They are classified into two groups. Group 1 analyzes the data from Part 1, while Group 2 is responsible for the data of Part 2. Members in each group work independently and then align conflicting results. Each group focuses only on their respective portion of the data, without interfering with each other.
Figure 2: The process of data pre-processing.
3.2.1 Filtering Extraneous Data
@@
In this step, we primarily filter out results unrelated to malicious code. Group 1 removes irrelevant data of Part 1, including empty code repositories, advertisements without substantial content, and web pages with only titles. Additionally, duplicate results and results unrelated to malicious code, such as evaluations or fixes using LLM for malicious code or security vulnerabilities of LLM itself, also be removed. Group 2 primarily selects high-quality malicious code repositories from Part 2, excluding repositories unrelated to the topic or invalid repositories, such as malicious code detection tools or repositories that do not provide access to source code directly. Furthermore, to ensure an adequate number of collected malicious code instances, repositories with at least $200$ stars will be considered [50]. After alignment within the groups, Group 1 obtains a total of $306$ results, including $128$ literature references, five code repositories, and $173$ relevant web pages and articles. Group 2 collects $93$ code repositories.
3.2.2 Extracting Malicious Code Instances
@@
In this step, both groups are tasked with obtaining instances of malicious code generated by LLMs from the filtered results obtained in the previous stage. This forms the foundation of our LLM malicious code prompt dataset. Group 1 extracts descriptions of malicious code generated by LLMs from literature, news articles, and code repositories. They also identify and analyze malicious code snippets to determine their functionality. Group 2, on the other hand, selects usable malicious code functionality from the malicious code repositories. They exclude incomplete or ambiguous code and functions that are unrelated to the repository description or purpose, such as data visualization or graphical user interfaces. Additionally, to maintain the quality of collected instances, Group 2 members will exclude functions with fewer than five lines of code. After alignment within the groups, Group 1 and Group 2 have selected $129$ and $395$ instances of malicious code, respectively.
@@
In summary, we collect a total of $524$ malicious code instances, encompassing both LLM-generated samples and a comprehensive survey of prominent open-source malicious code. thereby encompassing potential scenarios for malicious code generation.
3.3 The Construction Process of Prompt Dataset
@@
In this process, we construct a prompt dataset called MCGTest for LLM malicious code generation, based on the data collected from Section 3.2.
Figure 3: The construction process of MCGTest.
Fig. 3 illustrates the construction process of MCGTest. This process contains three steps. First, we create functional summaries of the malicious code instances extracted in Section 3.2. Next, we filter these instances to exclude those that do not meet our requirements, such as redundant functionalities. Finally, we use the filtered set of malicious code functionalities to construct a dataset of prompts for malicious code generation.
3.3.1 Summary of Malicious Code Functionality
@@
In this step, we aim to collect comprehensive information about malicious code’s functionality, including the overall malicious purpose of each instance and the specific malicious functions within them. To achieve this, we establish a Malicious Function Set (MFS) to compile all these functionalities. First, we summarize the functionality of each collected malicious code instance based on descriptions from their respective repositories or literature and add these summaries to the MFS. Next, we divide all instances into individual functions and use GPT-4 [51], an advanced LLM, to generate code summaries for these functions. Finally, we incorporate these summaries into the MFS. At this point, the MFS contains both the overall malicious purposes of all instances and the specific malicious functions found within them, ensuring a thorough representation of malicious functionalities at both the instance and function levels within our dataset.
3.3.2 Filtering Malicious Functionalities
@@
In this step, we filter malicious functionalities, since not all of them are malicious in MFS. For example, some serve as copyright declarations. Additionally, there may be redundancy in functionalities across different functions. Therefore, we need to filter out malicious code functionalities from MFS. During the filtering process, we employ a closed card sorting method to ensure the accuracy of results. Closed card sorting is one of the most efficient methods for organizing information into logical groups [52]. Two participants are involved; both have over four years of programming experience. The card title is code functionalities, and the description consists of code or descriptions from the code repository/literature. Both participants read and filter cards according to specified criteria, aligning their results. The overall filtering rules are as follows:
@@
3.3.3 Creating the Malicious Prompt Dataset
@@
In this step, we construct the MCGTest dataset based on $406$ filtered cards. We primarily use the titles on each card describing the malicious code functionality to craft corresponding prompts. As a result, MCGTest comprises $406$ prompts. The prompt creation task involved three participants, each with an average of over four years of programming experience. Initially, each prompt was assigned to two participants: one for writing and one for review. In cases of disagreement, the third participant would facilitate a discussion to reach a consensus. For the prompt format, we drew inspiration from MBPP and HumanEval [53], two well-known datasets for evaluating LLM code generation performance. Each prompt consists of a function description and a function name, as illustrated in Fig. 4. MCGTest is designed to be compatible with various LLMs. Typically, an LLM has multiple versions, differing mainly in parameter count and whether they are Base or Chat versions [27, 54]. Base versions are usually continuation models with limited human interaction capabilities, while Chat versions can engage in dialogue. MCGTest’s prompt format is compatible with both versions.
@@
In summary, we construct $406$ malicious code generation tasks for MCGTest. These tasks include real instances of LLM-generated malicious code as well as prominent potential scenarios. Additionally, MCGTest is compatible with LLMs and can be easily accessed from the open-source repository referenced in this paper. Beyond designing our watermarking scheme based on MCGTest, it will inspire further research on LLM-generated malicious code.
Figure 4: The prompt format of MCGTest.
4 MCGMark:An Encodable and Robust online watermark for LLM code generation
In this section, we introduce MCGMark, a method for embedding encodable watermarks over LLM code generation. MCGMark allows us to trace the developer of the generated malicious code. The watermark is implicit, and its robustness is enhanced by leveraging the structural features of the malicious code, making it difficult to be aware and resistant to tampering.
Figure 5: The overview of watermark embedding and detection.
4.1 Overview
Fig. 5 outlines the watermark embedding and detection process of MCGMark. Each process comprises five steps. It is noteworthy that both the embedding and detection of the watermark do not require the introduction of additional models or external data. Furthermore, during the watermark detection phase, there is no need to reload the LLM. This aligns well with the practical needs of SSPs.
@@ 4.1 Overview
For the watermark embedding process. MCGMark first initializes the watermark based on the user’s ID. The watermark consists of detection bits and error-correction bits, which collectively represent the user’s ID (Section 4.3.1 and Section 4.5.1). Subsequently, MCGMark partitions the LLM’s vocabulary and embeds multi-encoded watermarks by controlling the LLM’s token selection (Section 4.2). To mitigate the impact on code generation quality during watermark embedding, MCGMark then processes probability outliers in the vocabulary to ensure the LLM selects high-probability tokens, and updates the error-correction bit information (Section 4.3). Next, to enhance watermark robustness, we design a watermark skipping strategy for MCGMark based on code structure and syntax. MCGMark implements this strategy based on the generated code elements, ensuring that watermarks are not added to easily modifiable code elements. Finally, as code generation progresses, watermarks are embedded in a round-robin fashion to further enhance their robustness (Section 4.4).
@@
For the watermark detection process. MCGMark first tokenizes the code into a sequence of tokens with the tokenizer. Next, MCGMark removes tokens from the sequence that would have been skipped during the embedding process, based on the employed skip strategy. Subsequently, MCGMark partitions the vocabulary and examines the vocabulary membership of tokens in the sequence to recover the watermark information. Since the watermark is embedded in multiple rounds, MCGMark then trims the recovered watermark. Finally, MCGMark reconstructs the user’s ID from the multiple segments of the trimmed watermark (Section 4.5.2).
4.2 Encodable Watermark Embedding
@@ 4.2.1 LLM Code Generation Process
LLM Code Generation Process serves as the foundation for watermark embedding in MCGMark. In a typical LLM code generation process, the LLM maintains a token-level vocabulary $V=\left\{v_{0},v_{1},\cdots,v_{n}\right\}$, typically comprising approximately $3.2\times 10^{4}$ tokens (e.g., the DeepSeek-Coder-6.7b utilizes a vocabulary of $32,022$ tokens) [55]. When a prompt $R$ is input into the model $M$, it first employs a tokenizer to segment $R$ into token-level components, $R\Rightarrow T=\left\{t_{0},t_{1},\cdots,t_{m}\right\}$. Subsequently, $M$ computes the probability distribution $P_{1}=\left\{P_{0}^{\prime},P_{1}^{\prime},\cdots,P_{n}^{\prime}\right\}$ for all tokens in the vocabulary based on $T$. Then, $M$ selects the highest probability token $v_{i}^{\prime}(0\leqslant i\leqslant n)$ from the vocabulary as the generated token and appends it to the prompt $R$. This process transforms the prompt $R$ into $R^{\prime}=\left\{t_{0},t_{1},\cdots,t_{m},v_{i}^{\prime}\right\}$. $R^{\prime}$ is then fed back into model $M$, and this process iterates until a predetermined generation length $L$ is reached. The final output of the model is represented as $R^{(L)}=\left\{t_{0},t_{1},\cdots,t_{m},v_{i}^{\prime},v_{i}^{\prime\prime},\cdots,v_{i}^{(L)}\right\}$, where $\left\{v_{i}^{\prime},v_{i}^{\prime\prime},\cdots,v_{i}^{(L)}\right\}$ constitutes the generated code [56].
4.2.2 Watermark Embedding
@@
This step delineates the watermark embedding process in MCGMark. Inspired by the study [25], MCGMark incorporates watermarking by modifying the vocabulary $V=\left\{v_{0},v_{1},\cdots,v_{n}\right\}$. MCGMark first generates a random set $D$ ($0<|D|\leqslant|V|$) with a hash value $H$. The elements in $D$ are unique, increasing integer values representing positions in the vocabulary. These elements form the selected vocabulary $\mathbb{A}$, while the remaining elements constitute vocabulary $\mathbb{B}$. MCGMark controls the LLM’s token selection by modifying the probability distribution of the vocabulary, constraining it to choose from the selected vocabulary. Simultaneously, to prevent excessive dependence on $H$ in vocabulary partitioning, MCGMark applies a pseudo-random augmentation to $H$. This approach ensures that $H$ remains variable while maintaining it reproducible. As all generated tokens stem from randomly chosen vocabulary subsets, the LLM-generated code inherently differs from manually written code. The probability of complete overlap between manually-generated and model-generated code is merely $\frac{1}{2^{L}}$ [25]. This low probability ensures the effectiveness of the code watermark.
4.2.3 Encoding Watermark
@@
MCGMark encodes the watermark information to represent user-specific data. Specifically, MCGMark achieves encoding watermark embedding by modifying the probabilities of the LLM vocabulary. For embedding watermark $w_{w}$ into token $v_{i}^{(v)}$, MCGMark controls the LLM’s selection as follows:
@@
To ensure watermark correctness, MCGMark must guarantee that the model selects elements from the predefined vocabulary. This requires that the token with the highest probability in the modified vocabulary is present in the selected vocabulary. After generating element $v_{i}^{(v-1)}$, MCGMark obtains the probabilities of all tokens in the current vocabulary and calculates: $P_{\text{gap }}=\max\left\{P_{v}\right\}-\min\left\{P_{v}\right\}~{}.$
MCGMark then adds $P_{\text{gap}}$ to all elements in the selected vocabulary to ensure this vocabulary contains the highest probability token.
Thus, MCGMark completes the online multi-encoding watermark embedding in code generation.
4.3 Ensuring Watermarked Code Quality
In this process, MCG mitigates the impact of watermarking on code quality by addressing probability outliers in the vocabulary.
4.3.1 Code Quality Enhancement with Outliers
To minimize the impact of watermarks on LLM code generation quality and process high-probability tokens of the vocabulary, MCGMark develops a watermark quality enhancement algorithm based on the probability outlier values of the LLM’s vocabulary, as shown in Algorithm 1. This algorithm addresses potential issues arising from vocabulary partitioning, which may lead to incorrect selection of deterministic tokens and propagate errors in subsequent code generation. Leveraging the powerful generation capability of LLM, MCGMark implements the following strategy to ensure code generation quality during watermark embedding.
  1. 1.

### v2

@@ 3.2. Data Pre-processing
To identify both real instances and potential scenarios of LLM-generated malicious code, we design a two-step data preprocessing pipeline, as illustrated in Fig. 2. First, we remove irrelevant data, such as empty repositories. Then, we extract representative malicious code instances from the remaining data to analyze their functionality and construct the MCGTest dataset.
@@
To ensure the reliability of this analysis. Four researchers with over four years of software development experience are assigned to filter and analyze the results. They are classified into two groups. Group 1 analyzes the data from Part 1, while Group 2 is responsible for Part 2. Members in each group work independently and then align conflicting results. Each group focuses only on their respective portion of the data, without interfering with each other.
Figure 2. The process of data pre-processing.
@@
Filtering Extraneous Data. In this step, we primarily filter out results unrelated to malicious code. Group 1 removes irrelevant data of Part 1, including empty code repositories, advertisements without substantial content, and web pages with only titles. Additionally, duplicate results and results unrelated to malicious code, such as evaluations or fixes using LLM for malicious code or security vulnerabilities of LLM itself, also be removed. Group 2 primarily selects high-quality malicious code repositories from Part 2, excluding repositories unrelated to the topic or invalid repositories, such as malicious code detection tools or repositories that do not provide access to source code directly. Furthermore, to ensure an adequate number of collected malicious code instances, repositories with at least $200$ stars will be considered (Sülün et al., 2024). After alignment within the groups, Group 1 obtains a total of $306$ results, including $128$ literature references, five code repositories, and $173$ relevant web pages and articles. Group 2 collects $93$ code repositories.
@@
Extracting Malicious Code Instances. In this step, both groups are tasked with obtaining instances of malicious code generated by LLMs from the filtered results obtained in the previous stage. This forms the foundation of our LLM malicious code prompt dataset. Group 1 extracts descriptions of malicious code generated by LLMs from literature, news articles, and code repositories. They also identify and analyze malicious code snippets to determine their functionality. Group 2, on the other hand, selects usable malicious code functionality from the malicious code repositories. They exclude incomplete or ambiguous code and functions that are unrelated to the repository description or purpose, such as data visualization or graphical user interfaces. Additionally, to maintain the quality of collected instances, Group 2 members will exclude functions with fewer than five lines of code. After alignment within the groups, Group 1 and Group 2 have selected $129$ and $395$ instances of malicious code, respectively.
@@
In summary, we collected 524 malicious code instances, including LLM-generated samples and key open-source malicious code, covering potential generation scenarios.
3.3. The Construction Process of Prompt Dataset
@@
In this process, we construct the MCGTest prompt dataset for LLM-based malicious code generation using data collected in Section 3.2. As shown in Fig. 3, the process involves three steps: (1) summarizing the functionality of malicious code instances; (2) filtering out redundant or unsuitable cases; and (3) creating prompts based on the remaining instances.
@@
Summary of Malicious Code Functionality. In this step, we aim to collect comprehensive information about malicious code’s functionality, including both the overall intent of each instance and its specific malicious components. To achieve this, we establish a Malicious Function Set (MFS) to consolidates these functionalities. First, we summarize the functionality of each malicious code instance based on descriptions from their repositories or literature and add these summaries to the MFS. Next, we divide all instances into individual functions and use GPT-4 (Maniparambil et al., 2023), an advanced LLM, to generate code summaries for these functions. Finally, we incorporate these summaries into the MFS. As a result, the MFS captures malicious behavior at both the instance and function levels, providing a comprehensive view of malicious code functionalities within our dataset.
@@
Filtering Malicious Functionalities. In this step, we filter malicious functionalities because not all of them are necessary in MFS. For example, copyright declarations or redundant functionalities across different functions. To ensure filtering accuracy, we employ a closed card sorting method to ensure the accuracy of results. Closed card sorting is one of the most efficient methods for organizing information into logical groups (Chen et al., 2022). Two participants are involved, both have over four years of programming experience. The card title is code functionalities, and the description consists of code or descriptions from the code repository/literature. Both participants read and filter cards according to specified criteria, aligning their results. The overall filtering rules are as follows:
@@
Figure 3. The construction process and the prompt format of MCGTest.
@@
Creating the Malicious Prompt Dataset. In this step, we construct the MCGTest dataset, which comprises 406 prompts derived from filtered cards describing malicious code functionality. Three participants, each with over four years of programming experience, collaborated on creating these prompts. Every prompt was initially drafted and reviewed by two participants, with a third participant resolving any disagreements. For the prompt format, we drew inspiration from MBPP and HumanEval (Zheng et al., 2023), two well-known datasets for evaluating LLM code generation performance. Each prompt consists of a function description and a function name, as illustrated in Fig. 3. MCGTest is designed to be compatible with various LLMs. Typically, an LLM has multiple versions, differing mainly in parameter count and whether they are Base or Chat versions (He and Vechev, 2023; Chang et al., 2023). Base versions are usually continuation models with limited human interaction capabilities, while Chat versions can engage in dialogue. MCGTest’s prompt format is compatible with both versions.
@@
In summary, we construct $406$ malicious code generation tasks for MCGTest. These tasks include real instances of LLM-generated malicious code as well as prominent potential scenarios.
4. MCGMark:An Encodable and Robust watermark for LLM code generation
In this section, we introduce MCGMark, a method for embedding encodable watermarks during LLM code generation.
4.1. Overview
Fig. 4 outlines the watermark embedding and detection process of MCGMark. Each process comprises five steps.
@@ 4.1. Overview
Watermark embedding process. MCGMark first initializes the watermark based on the user’s ID. The watermark consists of detection bits and error-correction bits, which collectively represent the user’s ID (Section 4.3.1 and Section 4.5.1). Subsequently, MCGMark partitions the LLM’s vocabulary and embeds multi-encoded watermarks by controlling the LLM’s token selection (Section 4.2). To mitigate the impact on code generation quality during watermark embedding, MCGMark then processes probability outliers in the vocabulary to ensure the LLM selects high-probability tokens, and updates the error-correction bit information (Section 4.3). Next, to enhance watermark robustness, we design a watermark skipping strategy for MCGMark based on code structure and syntax. MCGMark implements this strategy based on the generated code elements, ensuring that watermarks are not added to easily modifiable code elements. Finally, as code generation progresses, watermarks are embedded in a round-robin fashion to further enhance their robustness (Section 4.4).
It is important to note that MCGMark operates independently of the LLM’s internal generation process. It does not interrupt or roll back token generation, nor does it rely on any additional models or external databases during embedding. Moreover, MCGMark is fully decoupled from the LLM architecture and does not participate in its neural computations. As a result, techniques such as fine-tuning, distillation, or prompt engineering have no effect on the watermarking process.
@@
Watermark detection process. MCGMark first tokenizes the code into a sequence of tokens with the tokenizer. Next, MCGMark removes tokens that would have been skipped during the embedding process, based on the employed skip strategy. Subsequently, MCGMark partitions the vocabulary and examines the vocabulary membership of tokens in the sequence to recover the watermark information. Since the watermark is embedded in multiple rounds, MCGMark then trims the recovered watermark. Finally, MCGMark reconstructs the user’s ID from the multiple segments of the trimmed watermark (Section 4.5.2).
It is important to note that MCGMark does not require simulating the LLM’s code generation process or accessing the LLM during detection. Only the tokenizer is required, which is used to split the code into tokens. This allows MCGMark to identify the vocabulary group each token belongs to and recover the embedded watermark information.
Figure 4. The overview of watermark embedding and detection.
4.2. Encodable Watermark Embedding
@@
LLM Code Generation Process. In a typical LLM code generation process, the LLM maintains a token-level vocabulary $V=\left\{v_{0},v_{1},\cdots,v_{n}\right\}$, typically comprising approximately $3.2\times 10^{4}$ tokens (e.g., the DeepSeek-Coder-6.7b utilizes a vocabulary of $32,022$ tokens) (Kandpal et al., 2023). When a prompt $R$ is input into the model $M$, it first employs a tokenizer to segment $R$ into token-level components, $R\Rightarrow T=\left\{t_{0},t_{1},\cdots,t_{m}\right\}$. Subsequently, $M$ computes the probability distribution $P_{1}=\left\{P_{0}^{\prime},P_{1}^{\prime},\cdots,P_{n}^{\prime}\right\}$ for all tokens in the vocabulary based on $T$. Then, $M$ selects the highest probability token $v_{i}^{\prime}(0\leqslant i\leqslant n)$ from the vocabulary as the generated token and appends it to the prompt $R$. This process transforms the prompt $R$ into $R^{\prime}=\left\{t_{0},t_{1},\cdots,t_{m},v_{i}^{\prime}\right\}$. $R^{\prime}$ is then fed back into model $M$, and this process iterates until a predetermined generation length $L$ is reached. The final output of the model is represented as $R^{(L)}=\left\{t_{0},t_{1},\cdots,t_{m},v_{i}^{\prime},v_{i}^{\prime\prime},\cdots,v_{i}^{(L)}\right\}$, where $\left\{v_{i}^{\prime},v_{i}^{\prime\prime},\cdots,v_{i}^{(L)}\right\}$ constitutes the generated code (Li et al., 2023b).
Watermark Embedding. Inspired by the study (Kirchenbauer et al., 2023a), MCGMark encodes binary information by dividing the LLM’s vocabulary into two parts and sampling tokens from these parts based on the user’s ID. The vocabulary division is performed using a pseudo-random partitioning process on hash seeds, which is dynamically adjusted according to specific rules. This approach preserves the random characteristics of the vocabulary during embedding while allowing the randomness to be accurately reproduced during detection (Hu et al., 2023; Christ et al., 2024).
@@
MCGMark incorporates watermarking by modifying the vocabulary $V=\left\{v_{0},v_{1},\cdots,v_{n}\right\}$. MCGMark first generates a random set $D$ ($0<|D|\leqslant|V|$) with a hash value $H$. The elements in $D$ are unique, increasing integers representing vocabulary positions. They define the selected vocabulary $\mathbb{A}$, with the rest forming $\mathbb{B}$. MCGMark adjusts the LLM’s probability distribution to constrain token selection to $\mathbb{A}$. To reduce reliance on $H$ in partitioning, it applies a pseudo-random augmentation to $H$. This approach ensures that $H$ remains variable while maintaining it reproducible. As all generated tokens stem from randomly chosen vocabulary subsets, the LLM-generated code inherently differs from manually written code. The probability of complete overlap between manually-generated and model-generated code is merely $\frac{1}{2^{L}}$ (Kirchenbauer et al., 2023a). This low probability ensures the effectiveness of the code watermark.
@@
Encoding Watermark. MCGMark encodes the watermark to represent user-specific data. Specifically, MCGMark achieves encoding watermark embedding by modifying the probabilities of the LLM vocabulary. For embedding watermark $w_{w}$ into token $v_{i}^{(v)}$, MCGMark controls the LLM’s selection as follows:
  1. (1)
@@
To ensure watermark correctness, MCGMark guarantee that the model selects elements from the predefined vocabulary. This requires that the token with the highest probability in the modified vocabulary is present in the selected vocabulary. After generating element $v_{i}^{(v-1)}$, MCGMark obtains the probabilities of all tokens in the current vocabulary and calculates: $P_{\text{gap }}=\max\left\{P_{v}\right\}-\min\left\{P_{v}\right\}~{}.$ MCGMark then adds $P_{\text{gap}}$ to all elements in the selected vocabulary to ensure this vocabulary contains the highest probability token.
Thus, MCGMark completes encodable watermark embedding in code generation.
4.3. Ensuring Watermarked Code Quality
In this process, MCGMark maintains the quality of the generated code by guiding the LLM to select from high-quality token candidates.
1:  Input: Prompt: $R$, watermark: $W$, the number of new tokens: $L$, threshold for $P_{\text{dis }}$: $Thr\_P_{\text{dis }}$, a hash key $H$.
2:  for $l=0,1,\cdots,L$ do
3:     Obtain $V=\left\{v_{0},v_{1},\cdots,v_{n}\right\}$ and $P_{l}=\left\{P_{0}^{l},P_{1}^{l},\cdots,P_{n}^{l}\right\}$.
4:     Calculate the outliers of $P_{l}$.
5:     if $F_{\text{upper }}\neq\emptyset$ then
6:        Case 1: $|F_{\text{upper }}|=1$,$D^{{}^{\prime}}=D\cup F_{\text{upper }}$.
7:        Case 2: $|F_{\text{upper }}|\geq 2$, $D^{{}^{\prime}}=D\cup F_{\text{upper }}[0:\lceil\frac{F_{\text{upper }}}{2}\rceil]$ with $H$.
8:     end if
9:     Sample $v_{i}^{(l)}$ from $V$.
10:     if $v_{i}^{(l)}\in F_{\text{upper}}~{}\&~{}v_{i}^{(l)}\notin D$ then
11:        Setting the watermark’s error-correction bit to $1$.
12:     end if
13:  end for
Code Quality Enhancement with Outliers. To minimize the impact of watermarks on LLM code generation quality, MCGMark develops a watermark quality enhancement algorithm based on the probability outlier of the vocabulary, as shown in Algorithm 1. This algorithm addresses potential issues arising from vocabulary partitioning, which may lead to incorrect selection of deterministic tokens and propagate errors in subsequent code generation. Leveraging the powerful generation capability of LLM, MCGMark implements the following strategy to ensure code generation quality during watermark embedding.
  1. (1)